# SynSeg: Feature Synergy for Multi-Category Contrastive Learning in Open-Vocabulary Semantic Segmentation

### v1

@@
Abstract
@@
Semantic segmentation in open-vocabulary scenarios presents significant challenges due to the wide range and granularity of semantic categories. Existing weakly-supervised methods often rely on category-specific supervision and ill-suited feature construction methods for contrastive learning, leading to semantic misalignment and poor performance. In this work, we propose a novel weakly-supervised approach, SynSeg, to address the challenges. SynSeg performs Multi-Category Contrastive Learning (MCCL) as a stronger training signal with a new feature reconstruction framework named Feature Synergy Structure (FSS). Specifically, MCCL strategy robustly combines both intra- and inter-category alignment and separation in order to make the model learn the knowledge of correlations from different categories within the same image. Moreover, FSS reconstructs discriminative features for contrastive learning through prior fusion and semantic-activation-map enhancement, effectively avoiding the foreground bias introduced by the visual encoder. In general, SynSeg effectively improves the abilities in semantic localization and discrimination under weak supervision. Extensive experiments on benchmarks demonstrate that our method outperforms state-of-the-art (SOTA) performance. For instance, SynSeg achieves higher accuracy than SOTA baselines by 4.5% on VOC, 8.9% on Context, 2.6% on Object and 2.0% on City.
Introduction
@@
Semantic segmentation is a fundamental task in computer vision that focuses on the classification of each pixel in an image with respect to semantic categories. This task has numerous practical applications, including autonomous driving (Cakir et al. 2022; Wang et al. 2022a), medical image analysis (Kar, Nath, and Neog 2021; Jaus et al. 2025) and embodied intelligence (Çağrı Kaymak and Uçar 2018). However, due to the wide range and variability of object categories in the open vocabulary scenarios encountered in real-world tasks, traditional semantic segmentation methods with fixed categories are often insufficient.
@@
To overcome these limitations, many Open Vocabulary Semantic Segmentation (OVSS) approaches have been developed recently (Xu et al. 2022; Mukhoti et al. 2023; Cha, Mun, and Roh 2023; Lan et al. 2024; Liu et al. 2022; Xing et al. 2023; Wu et al. 2024; Ding, Wang, and Tu 2023). OVSS aims to segment any object category, including those not explicitly defined during training, enabling more flexible and scalable scene understanding. Training such OVSS systems typically requires large amounts of pre-annotated data at pixel level, while semantic annotation in open vocabulary scenarios is both costly and challenging (Cho et al. 2024; Ghiasi et al. 2022; Xu et al. 2023). Manual annotation leads to substantial human labor and is prone to the detrimental effects of poor-quality annotations. In contrast, weakly-supervised learning methods that incorporate semantic text cues into the semantic segmentation tasks via visual-text alignment techniques offer a compelling alternative.
@@
Nevertheless, existing weakly-supervised OVSS solutions often fail to achieve satisfactory performance due to the lack of accurate and concrete supervisory signals. GroupViT (Xu et al. 2022) and VilSeg (Liu et al. 2022) are the earliest approaches in the field of OVSS and have paved the way for numerous subsequent studies (Lai 2024; Luo et al. 2023; Ding, Wang, and Tu 2023; Yi et al. 2023). They introduce a simple image-text alignment architecture, which is shown in Fig. 1(a). Specifically, they operate by grouping local visual features and matching these clusters with text embeddings at test time to generate segmentation masks. However, during training, text embeddings are aligned with a global image representation rather than the detailed region-level features used during inference. This mismatch leads to a notable inconsistency between training and testing, which may impede the model’s ability to fully capture fine-grained semantic details.
(a) image-text alignment (Xu et al. 2022)
(b) region-text (Cha, Mun, and Roh 2023) / region-word alignment (Wu et al. 2024)
@@
Some other approaches like TCL (Cha, Mun, and Roh 2023) improved on the above by introducing region-level visual-text alignment training objective, which is shown in Fig. 1(b). These methods leverage the directional cues provided by textual descriptions to jointly guide the regions’ segmentation in both training and testing process (Wu et al. 2024; Mukhoti et al. 2023). These schemes work well in simple scenes with sparse targets. They, however, encounter a substantial limitation when applied to scenarios with dense targets which are common in real-word open vocabulary settings (Caesar, Uijlings, and Ferrari 2018; Lin et al. 2014). This limitation comes from the fact that each segmentation region usually aligns to just one semantic category. And the loss functions in these solutions are usually designed in a category-specific way, focusing on only intra-category alignment. After all, in real-world scenes where multiple categories appear in close spatial proximity or even visually overlap with each other, intra-category alignment alone is insufficient to achieve strong semantic discrimination. Without a coordinated interplay of both intra- and inter-category alignment and separation mechanisms, the model struggles to disambiguate between overlapping or visually similar regions. As a result, this leads to the first challenge: existing OVSS methods lack explicit modeling of inter-category correlations during training, which limits their ability to distinguish semantically different targets in one image.
Moreover, as shown in Fig. 1(a) and (b), previous works typically rely on frozen, pre-trained vision encoders such as CLIP (Radford et al. 2021) to extract features for contrastive learning after the decoder. However, these features often exhibit limited discriminability in background regions, as the encoder is inherently biased toward salient foreground objects (Rao et al. 2022; Ding, Wang, and Tu 2023). Therefore, segmentation regions contaminated by background noise may still yield highly similar representations to clean regions. This hinders the rapid decline of the contrastive loss and reduces the overall learning efficiency. As a result, this leads to the second challenge: existing methods lack the ability to reconstruct representations that are well-suited for contrastive learning in open-vocabulary semantic segmentation.
@@
Visual-Language Models
@@
Vision-language models aim to bridge the gap between visual and textual modalities, enabling a broad range of multi-modal tasks such as image-text retrieval (Li et al. 2021; Radford et al. 2021), image captioning (Li et al. 2022; Zhang et al. 2021), open-vocabulary detection and recognition (Tang et al. 2024; Wang et al. 2022b; Kim et al. 2022; Wang et al. 2024b), and multi-source semantic segmentation (Deng et al. 2023; Guo et al. 2024). Among these models, CLIP (Contrastive Language-Image Pretraining) (Radford et al. 2021) introduced a framework for joint visual and textual representation learning using contrastive learning on a large-scale dataset of image-text pairs. It aligns image and text embeddings in a shared feature space, enabling zero-shot tasks and fundamental abilities. BLIP (Bootstrapped Language-Image Pretraining) (Li et al. 2022) enhances vision-language models with a multi-modal encoder-decoder architecture and combined generative and contrastive objectives. Its bootstrapped self-learning improves performance across datasets, making it effective for tasks such as image captioning and vision-language understanding.
Weakly-Supervised Open-Vocabulary Semantic Segmentation
@@
In early OVSS works (Xu et al. 2022; Liu et al. 2022; Luo et al. 2023), textual information does not participate in the computation or generation of segmentation masks but only plays a role in the subsequent matching of mask proposals with candidate semantic labels. Multiple later works follow this structure and explore improvements in segmentation accuracy and semantic alignment precision (Chen et al. 2023; Cai et al. 2023; Lai 2024; Xing et al. 2023). However, with such structure, the visual segmentation output remains unchanged regardless of the provided textual prompt. Furthermore, in the architectures based on the classic GroupViT (Xu et al. 2022), the number of output class token is fixed (e.g., 4 or 8), making it impossible to segment more semantic categories than this predefined limit. Since the text is not involved in the generation of mask proposals, and only a limited number of semantically irrelevant proposals are produced, this structure necessitates an additional step for global-level alignment, which in turn limits the overall learning efficiency. The inconsistency between region-level alignment during testing and global-level alignment during training also leads to poor performance in fine-grained segmentations.
@@
TCL (Cha, Mun, and Roh 2023) performs a different structure firstly to partially incorporate text labels into the mask generation process and apply region-text alignment during training. Typically, this occurs at the final step, where each pixel feature is compared with text label embeddings to cluster and assign it to a candidate label, determining which pixels are highly correlated with the semantic information to produce the final segmentation result. This paradigm is followed by later works (Mukhoti et al. 2023; Wang et al. 2024a; Sun et al. 2024). For example, CoDe (Wu et al. 2024) improves its work by introducing region-word alignment. In these works, the integration of textual information with the visual feature enables the mask proposal to refer to certain semantic more flexibly, thereby enhancing the granularity of segmentation alignment. However, This training approach mainly focuses on a single object or semantic category and does not exploit the relationships among different semantic objects within the same image.
@@
Furthermore, these works typically feed segmentation results into the CLIP models (Radford et al. 2021) again to construct visual features for loss computing. Since the CLIP model is inherently biased toward salient foreground objects, its visual representations tend to be sparse and less discriminative in background regions (Rao et al. 2022; Ding, Wang, and Tu 2023). In OVSS, when segmentation predictions inadvertently include background noise—pixels that actually belong to other semantic categories—such sparse representations hinder the model’s ability to produce more accurate and precise segmentation regions. Due to CLIP’s limited capacity for representing background content, the visual features of noisy segmentation regions often remain highly similar to those of cleaner regions, which serve as the intended training targets. This leads to optimization bottlenecks of the loss functions’ decline that weaken the effectiveness of contrastive learning, both in terms of intra- and, if applicable, inter-category alignment and separation.
Approach
@@
Figure 2: The pipeline of SynSeg. It illustrates the proposed Feature Synergy Structure and Multi-Category Contrastive Learning framework. During training, FiLM (Perez et al. 2018) fusion module, transformer decoder and the projector stay trainable, while the CLIP (Radford et al. 2021) encoders stay frozen. The projector is here to make sure the feature vectors in an appropriate dimension for later use.
Overview of SynSeg
Our pipeline, named SynSeg, is illustrated in Fig. 2. The pre-trained CLIP (Radford et al. 2021) visual and textual encoders process the input image and text labels, respectively, to extract single-modal embeddings. These embeddings are then fused using a Feature-wise Linear Modulation (FiLM) (Perez et al. 2018) module. Specifically, the FiLM module contains a small learnable MLP that takes the text features as input and outputs channel-wise scaling and shifting parameters. These parameters are then directly applied to the visual features to generate the conditional visual vectors. Meanwhile, textual embeddings are stored as semantic feature vectors for intra-category alignment later. During inference, the conditional visual features are passed through the decoder, generating semantic-activation maps, which are similar to class activation mapping (CAM) (Zhou et al. 2016). Then, the semantic-activation maps are thresholded to generate the final segmentation outputs.
Feature Synergy Structure
In the training phase, we introduce a new feature reconstruction framework, named Feature Synergy Structure (FSS), inspired by CCAM (Xie et al. 2022). Our focus is on the post-decoder reconstruction of features after the segmentation stage, and these reconstructed features are then fed into the subsequent contrastive learning. Rather than running the visual encoder a second time, we rely on FSS to perform this reconstruction.
The semantic-activation maps, which are class-specific heat maps, represent the continuous per-pixel response strength to a given semantic category and can be considered as attention maps for feature enhancement. Instead of applying thresholding, the semantic-activation maps are flattened and subsequently fused with conditional visual features, which are first projected to a compatible dimension via a convolutional-layer-based projector. This fusion is performed through matrix multiplication to generate the synergy feature vectors. By flattening and weighting with semantic activation maps, the synergy vector emphasizes spatial regions with high semantic relevance, preserving contextual cues.
Multi-Category Contrastive Learning
@@
Experiments
Experiment Setup
Training datasets. We use the public conceptual-12m (CC12M) (Changpinyo et al. 2021) as training dataset. After using the NLP functions from the SpaCy library (Honnibal and Montani 2017), nouns and noun phrases are extracted from these captions. Non-referential or irrelevant terms such as direction and unit nouns (e.g., southwest, pair, front) are filtered out, leaving meaningful text labels that provide weak supervision for the associated images. We resize the figures to the same pixel size of $224\times 224$ for training.
Evaluating datasets. To evaluate the open vocabulary semantic segmentation performance of our method, we test it on five commonly used challenging datasets: PASCAL VOC (VOC) (Everingham et al. 2010), Pascal Context (Context) (Mottaghi et al. 2014), City Scapes (City) (Cordts et al. 2016) , COCO Object (Object) and COCO Stuff (Stuff) (Lin et al. 2014; Caesar, Uijlings, and Ferrari 2018). Also, it should be noted that evaluations on VOC and Object treat unlabeled regions as an explicit category background, whereas those on Context, Stuff and City focus solely on labeled categories.
Baselines. To provide a comprehensive comparison, we select not only the latest but also classical weakly-supervised learning methods as baselines. The eight representative baselines are listed below: GroupViT (Xu et al. 2022), ViewCo (Ren et al. 2023), CoCu (Xing et al. 2023), OVSegmentor, TCL (Cha, Mun, and Roh 2023), CoDe (Wu et al. 2024), MGCA (Liu et al. 2025) and S-Seg(Lai 2024). All baseline methods are assumed to use a pre-trained ViT-B/16 vision backbone for fair comparison, except S-Seg (Lai 2024), which does not specify the type of vision model used in its paper. These methods may use extra training datasets such as CC3M (Sharma et al. 2018), YFCC14M (Thomee et al. 2016) and RedCaps12M (Desai et al. 2021).
Training Details. We use a pre-trained CLIP Vit-B/16 model as the encoder. The decoder follows the design of CLIPSeg (Lüddecke and Ecker 2022) and is initialized with the pre-trained weights from CLIPseg. For training, we keep the encoder frozen and only the decoder and the FiLM fusion module remain trainable. All of the experiments are performed on an RTX 4090 GPU.
Main Results
We compare our proposed method, SynSeg, against a range of classic and latest weakly-supervised open-vocabulary semantic segmentation approaches that rely on text annotations, across five widely-used benchmark datasets (Tab. 1). The table reports the mean Intersection-over-Union (mIoU) score for each method on individual datasets. The result in boldface achieves the best performance and the underlined result is the second best. SynSeg achieves the best overall performance, with an average mIoU of 38.7%, outperforming all baselines by a significant margin. Notably, our method sets a new state-of-the-art on four out of five datasets: VOC (62.2%), Context (41.8%), Object (34.9%), and City (30.9%).

### v2

# SynSeg: Feature Synergy for Multi-Category Contrastive Learning in End-to-End Open-Vocabulary Semantic Segmentation

@@
SynSeg: Feature Synergy for Multi-Category Contrastive Learning in End-to-End Open-Vocabulary Semantic Segmentation
Weichen Zhang1, Kebin Liu1, Fan Dang3, Zhui Zhu2, Xikai Sun2, Yunhao Liu1,2
1Global Innovation Exchange, Tsinghua University, Beijing, China
2Department of Automation, Tsinghua University, Beijing, China
3School of Software Engineering, Beijing Jiaotong University, Beijing, China
{weic_zhang23, z-zhu22, sxk23}@mails.tsinghua.edu.cn,
{kebinliu2021, yunhao}@tsinghua.edu.cn, dangfan@bjtu.edu.cn
Corresponding author.
@@
Semantic segmentation in open-vocabulary scenarios presents significant challenges due to the wide range and granularity of semantic categories. Existing weakly-supervised methods often rely on category-specific supervision and ill-suited feature construction methods for contrastive learning, leading to semantic misalignment and poor performance. In this work, we propose a novel weakly-supervised approach, SynSeg, to address the challenges. SynSeg performs Multi-Category Contrastive Learning (MCCL) as a stronger training signal with a new feature reconstruction framework named Feature Synergy Structure (FSS). Specifically, MCCL strategy robustly combines both intra- and inter-category alignment and separation in order to make the model learn the knowledge of correlations from different categories within the same image. Moreover, FSS reconstructs discriminative features for contrastive learning through prior fusion and semantic-activation-map enhancement, effectively avoiding the foreground bias introduced by the visual encoder. Furthermore, SynSeg is a lightweight end-to-end solution without using any mid-term output from large-scale pretrained models and capable for real-time inference. In general, SynSeg effectively improves the abilities in semantic localization and discrimination under weak supervision in an efficient manner. Extensive experiments on benchmarks demonstrate that our method outperforms state-of-the-art (SOTA) performance. Particularly, SynSeg achieves higher accuracy than SOTA baselines with a ratio from 6.9% up to 26.2%.
1 Introduction
@@
Semantic segmentation is a fundamental task in computer vision that focuses on the classification of each pixel in an image with respect to semantic categories. This task has numerous practical applications, including autonomous driving [3, 44], medical image analysis [18, 16] and embodied intelligence [56]. However, due to the wide range and variability of object categories in the open vocabulary scenarios encountered in real-world tasks, traditional semantic segmentation methods with fixed categories are often insufficient.
@@
To overcome these limitations, many Open Vocabulary Semantic Segmentation (OVSS) approaches have been developed recently [50, 32, 4, 23, 27, 49, 47, 11]. OVSS aims to segment any object category, including those not explicitly defined during training, enabling more flexible and scalable scene understanding. Training such OVSS systems typically requires large amounts of pre-annotated data at pixel level, while semantic annotation in open vocabulary scenarios is both costly and challenging [7, 13, 51]. Manual annotation leads to substantial human labor and is prone to the detrimental effects of poor-quality annotations. In contrast, weakly-supervised learning methods that incorporate semantic text cues into the semantic segmentation tasks via visual-text alignment techniques offer a compelling alternative.
@@
Nevertheless, existing weakly-supervised OVSS solutions often fail to achieve satisfactory performance due to the lack of accurate and concrete supervisory signals. GroupViT [50] and VilSeg [27] are the earliest approaches in the field of OVSS and have paved the way for numerous subsequent studies [22, 30, 11, 52]. They introduce a simple image-text alignment architecture, which is shown in Fig. 1(a). Specifically, they operate by grouping local visual features and matching these clusters with text embeddings at test time to generate segmentation masks. However, during training, text embeddings are aligned with a global image representation rather than the detailed region-level features used during inference. This mismatch leads to a notable inconsistency between training and testing, which may impede the model’s ability to fully capture fine-grained semantic details.
(a) image-text alignment [50]
(b) region-text [4] / region-word alignment [47]
@@
Some other approaches like TCL [4] improved on the above by introducing region-level visual-text alignment training objective, which is shown in Fig. 1(b). These methods leverage the directional cues provided by textual descriptions to jointly guide the regions’ segmentation in both training and testing process [47, 32]. These schemes work well in simple scenes with sparse targets. They, however, encounter a substantial limitation when applied to scenarios with dense targets which are common in real-word open vocabulary settings [1, 26]. This limitation comes from the fact that each segmentation region usually aligns to just one semantic category. And the loss functions in these solutions are usually designed in a category-specific way, focusing on only intra-category alignment. After all, in real-world scenes where multiple categories appear in close spatial proximity or even visually overlap with each other, intra-category alignment alone is insufficient to achieve strong semantic discrimination. Without a coordinated interplay of both intra- and inter-category alignment and separation mechanisms, the model struggles to disambiguate between overlapping or visually similar regions. As a result, this leads to the first challenge: existing OVSS methods lack explicit modeling of inter-category correlations during training, which limits their ability to distinguish semantically different targets in one image.
Moreover, as shown in Fig. 1(a) and (b), previous works typically rely on frozen, pre-trained vision encoders such as CLIP [34] to extract features for contrastive learning after the decoder. However, these features often exhibit limited discriminability in background regions, as the encoder is inherently biased toward salient foreground objects [35, 11]. Therefore, segmentation regions contaminated by background noise may still yield highly similar representations to clean regions. This hinders the rapid decline of the contrastive loss and reduces the overall learning efficiency. As a result, this leads to the second challenge: existing methods lack the ability to reconstruct representations that are well-suited for contrastive learning in open-vocabulary semantic segmentation.
@@ 2 Related Work
2.1 Visual-Language Models
@@
Vision-language models aim to bridge the gap between visual and textual modalities, enabling a broad range of multi-modal tasks such as image-text retrieval [24, 34], image captioning [25, 53], open-vocabulary detection and recognition [41, 45, 20, 46], and multi-source semantic segmentation [9, 14]. Among these models, CLIP (Contrastive Language-Image Pretraining) [34] introduced a framework for joint visual and textual representation learning using contrastive learning on a large-scale dataset of image-text pairs. It aligns image and text embeddings in a shared feature space, enabling zero-shot tasks and fundamental abilities. BLIP (Bootstrapped Language-Image Pretraining) [25] enhances vision-language models with a multi-modal encoder-decoder architecture and combined generative and contrastive objectives. Its bootstrapped self-learning improves performance across datasets, making it effective for tasks such as image captioning and vision-language understanding.
2.2 Weakly-Supervised OVSS Methods
Figure 2: The pipeline of SynSeg. It illustrates the proposed Feature Synergy Structure and Multi-Category Contrastive Learning framework. During training, FiLM [33] fusion module, transformer decoder and the projector stay trainable, while the CLIP [34] encoders stay frozen. The projector is here to make sure the feature vectors in an appropriate dimension for later use.
@@ 2.2 Weakly-Supervised OVSS Methods
In early OVSS works [50, 27, 30], textual information does not participate in the computation or generation of segmentation masks but only plays a role in the subsequent matching of mask proposals with candidate semantic labels. Multiple later works follow this structure and explore improvements in segmentation accuracy and semantic alignment precision [6, 2, 22, 49]. However, with such structure, the visual segmentation output remains unchanged regardless of the provided textual prompt. Furthermore, in the architectures based on the classic GroupViT [50], the number of output class token is fixed (e.g., 4 or 8), making it impossible to segment more semantic categories than this predefined limit. Since the text is not involved in the generation of mask proposals, and only a limited number of semantically irrelevant proposals are produced, this structure necessitates an additional step for global-level alignment, which in turn limits the overall learning efficiency. The inconsistency between region-level alignment during testing and global-level alignment during training also leads to poor performance in fine-grained segmentations.
@@
TCL [4] performs a different structure firstly to partially incorporate text labels into the mask generation process and apply region-text alignment during training. Typically, this occurs at the final step, where each pixel feature is compared with text label embeddings to cluster and assign it to a candidate label, determining which pixels are highly correlated with the semantic information to produce the final segmentation result. This paradigm is followed by later works [32, 43, 40]. For example, CoDe [47] improves its work by introducing region-word alignment. In these works, the integration of textual information with the visual feature enables the mask proposal to refer to certain semantic more flexibly, thereby enhancing the granularity of segmentation alignment. However, This training approach mainly focuses on a single object or semantic category and does not exploit the relationships among different semantic objects within the same image.
@@
Furthermore, these works typically feed segmentation results into the CLIP models [34] again to construct visual features for loss computing. Since the CLIP model is inherently biased toward salient foreground objects, its visual representations tend to be sparse and less discriminative in background regions [35, 11]. In OVSS, when segmentation predictions inadvertently include background noise—pixels that actually belong to other semantic categories—such sparse representations hinder the model’s ability to produce more accurate and precise segmentation regions. Due to CLIP’s limited capacity for representing background content, the visual features of noisy segmentation regions often remain highly similar to those of cleaner regions, which serve as the intended training targets. This leads to optimization bottlenecks of the loss functions’ decline that weaken the effectiveness of contrastive learning, both in terms of intra- and, if applicable, inter-category alignment and separation.
3 OVSS Based on Large Pretrained Models
Beyond the above weakly-supervised OVSS approaches, some works follow the training-free or inference-only paradigm [23, 39, 19]. These methods often depend on the powerful capabilities of the large-scale pretrained models by refining their mid-term outputs instead of end-to-end inference. Some other weakly supervised methods also fall under this paradigm [22, 54]. For example, ProxyCLIP [23] builds upon powerful pretrained segmentation models such as SAM [21] and DINO [17], while DPSeg [54] relies on Stable Diffusion [37] for visual prompting. We acknowledge the value of these explorations that leverage large pretrained models, however, they inevitably sacrifice lightweightness and efficiency.
To keep efficiency, we construct our method, SynSeg, in an end-to-end scheme without relying on any large pretrained models. We provide a simple comparison below in Table 1 to show the difference. The results are tested on an RTX 4090 GPU.
Method End-to-End Param. Lantency FPS
SynSeg (Ours) Yes 151M 14 ms 71
TCL Yes 178M 13 ms 76
ProxyCLIP No 243M 58 ms 17
DPSeg No $\sim$1.2B - -
Table 1: Comparison of model parameters and inference speed on different types of OVSS pipelines.
4 Approach
@@
4.1 Overview of SynSeg
Our pipeline, named SynSeg, is illustrated in Fig. 2. The pre-trained CLIP [34] visual and textual encoders process the input image and text labels, respectively, to extract single-modal embeddings. These embeddings are then fused using a Feature-wise Linear Modulation (FiLM) [33] module. Specifically, the FiLM module contains a small learnable MLP that takes the text features as input and outputs channel-wise scaling and shifting parameters. These parameters are then directly applied to the visual features to generate the conditional visual vectors. Meanwhile, textual embeddings are stored as semantic feature vectors for intra-category alignment later.
During inference, the conditional visual features are passed through the decoder, generating semantic-activation maps, which are similar to class activation mapping (CAM) [55]. Then, the semantic-activation maps are thresholded to generate the final segmentation outputs. Besides CLIP encoders, SynSeg does not utilize any large pretrained models and follows an end-to-end paradigm, which enables efficient inference.
4.2 Multi-Category Contrastive Learning
@@
4.3 Feature Synergy Structure
In the training phase, we introduce a new feature reconstruction framework, named Feature Synergy Structure (FSS), inspired by CCAM [48]. Our focus is on the post-decoder reconstruction of features after the segmentation stage, and these reconstructed features are then fed into MCCL. In the other words, FSS generate background and foreground feature vectors for each semantic category in one pass which are suitable for the subsequent MCCL.
Rather than running the visual encoder a second time, we rely on FSS to perform this reconstruction.The semantic-activation maps, which are class-specific heat maps, represent the continuous per-pixel response strength to a given semantic category and can be considered as attention maps for feature enhancement. Instead of applying thresholding, the semantic-activation maps are flattened and subsequently fused with conditional visual features, which are first projected to a compatible dimension via a convolutional-layer-based projector. This fusion is performed through matrix multiplication to generate the synergy feature vectors. By flattening and weighting with semantic activation maps, the synergy vector emphasizes spatial regions with high semantic relevance, preserving contextual cues.
5 Experiments
@@
5.1 Experiment Setup
Training datasets. We use the public conceptual-12m (CC12M) [5] as training dataset. After using the NLP functions from the SpaCy library [15], nouns and noun phrases are extracted from these captions. Non-referential or irrelevant terms such as direction and unit nouns (e.g., southwest, pair, front) are filtered out, leaving meaningful text labels that provide weak supervision for the associated images. We resize the figures to the same pixel size of $224\times 224$ for training.
Evaluating datasets. To evaluate the open vocabulary semantic segmentation performance of our method, we test it on five commonly used challenging datasets: PASCAL VOC (VOC) [12], Pascal Context (Context) [31], City Scapes (City) [8] , COCO Object (Object) and COCO Stuff (Stuff) [26, 1]. Also, it should be noted that evaluations on VOC and Object treat unlabeled regions as an explicit category background, whereas those on Context, Stuff and City focus solely on labeled categories.
Baselines. To provide a comprehensive comparison, we select not only the latest but also classical weakly-supervised learning methods as baselines. The eight representative baselines are listed below: GroupViT [50], ViewCo [36], CoCu [49], OVSegmentor, TCL [4], CoDe [47], MGCA [28] and S-Seg[22]. All baseline methods are assumed to use a pre-trained ViT-B/16 vision backbone for fair comparison, except S-Seg [22], which does not specify the type of vision model used in the inference stage. These methods may use extra training datasets such as CC3M [38], YFCC14M [42] and RedCaps12M [10].
Training Details. We use a pre-trained CLIP Vit-B/16 model as the encoder. The decoder follows the design of CLIPSeg [29] and is initialized with the pre-trained weights from CLIPseg. For training, we keep the encoder frozen and only the decoder and the FiLM fusion module remain trainable. All of the experiments are performed on an RTX 4090 GPU.
5.2 Main Results
We compare our proposed method, SynSeg, an end-to-end lightweight solution against a range of classic and latest weakly-supervised open-vocabulary semantic segmentation approaches that rely on text annotations, across five widely-used benchmark datasets (Tab. 2). The table reports the mean Intersection-over-Union (mIoU) score for each method on individual datasets. The result in boldface achieves the best performance and the underlined result is the second best. SynSeg achieves the best overall performance, with an average mIoU of 38.7%, outperforming all baselines by a significant margin. Notably, our method sets a new state-of-the-art on four out of five datasets: VOC (62.2%), Context (41.8%), Object (34.9%), and City (30.9%). To better understand, SynSeg achieves higher segmentation accuracy than SOTA baselines with a ratio from 6.9% up to 26.2%.